A flask weighs 50.000 g empty and E g when filled with a gas vapor at 100C and 758 mm Hg.

Taking the volume of the flask to be 225.0 mL, calculate the molar mass of the gas.
Answer #1

Mass of gas = total mass - flask mass = (E - 50) grams

Let n is the number of moles of gas.

T = 100° C = 100+273 = 373 K

R = 0.0821 L atm/mol K

P = 758 mmHg = 758/760 = 0.9974 atm

V = 225 ml = 0.225 liters

Using ideal gas equation;

PV = nRT

n = PV/RT = (0.9974 * 0.225) / (0.0821*373) = 0.01 moles

Moles = mass/molar mass

Molar mass = mass/moles = (E - 50)/0.01 = 100 * (E -50) g/mol ...Answer

Let me know if any doubts.